Which is the sum of the infinite progression 3/2, 1, 2/3, 4/9 ......?
Please put the solution, if it's possible.. thanks in advance

Answer by J2R2R(94)

What is the sum of the infinite progression 3/2, 1, 2/3, 4/9 ......?

This is a geometric progression and the sum to infinity for geometric progressions when the magnitude of the ratio is less than 1 is a/(1-r), where a is the first term and r is the ratio.

In this example, a = 3/2, r = 2/3, therefore 1-r = 1/3

Therefore, the sum to infinity = (3/2)/(1/3) = (3 x 3)/(2 x 1) = 9/2 = 4.5